# Free3D: Consistent Novel View Synthesis without 3D Representation

### v1

@@ Abstract
We introduce Free3D, a simple approach designed for open-set novel view synthesis (NVS) from a single image. Similar to Zero-1-to-3, we start from a pre-trained 2D image generator for generalization, and fine-tune it for NVS. Compared to recent and concurrent works, we obtain significant improvements without resorting to an explicit 3D representation, which is slow and memory-consuming or training an additional 3D network. We do so by encoding better the target camera pose via a new per-pixel ray conditioning normalization (RCN) layer. The latter injects pose information in the underlying 2D image generator by telling each pixel its specific viewing direction. We also improve multi-view consistency via a light-weight multi-view attention layer and multi-view noise sharing. We train Free3D on the Objaverse dataset and demonstrate excellent generalization to various new categories in several new datasets, including OminiObject3D and GSO. We hope our simple and effective approach will serve as a solid baseline and help future research in NVS with more accuracy pose. The project page is available at https://chuanxiaz.com/free3d/.
@@
Figure 1: Free3D novel view synthesis in an open-set setting. Given a single input view, our approach synthesizes consistent $360^{\circ}$ videos accurately and with no explicit 3D representation. Trained on Objaverse only, it generalizes well to new datasets and categories.
@@ 1 Introduction
Novel view synthesis (NVS) has developed rapidly in the recent past [5, 49, 6, 11, 38], fuelled in part by NeRF [48]. However, many such NVS methods require to optimize a new model from scratch for each scene, and are thus impractical in many applications. Inspired by generative models [39, 25, 67], other NVS methods generalize beyond a single scene to a specific object category, or even to unstructured datasets of objects [50, 34, 85, 32, 76, 47, 69]. Remarkably, these methods, by utilizing the data prior, can generate new views starting from a single image. However, they trade off some quality in order to achieve generality, which is still insufficient to build a 3D foundation model. Going beyond, in this work we consider the NVS problem in an open-set setting, generalizing beyond a single object, category, and dataset, yet still utilizing a single image as input, and aiming at obtaining high-quality outputs.
@@
There are two primary quality targets: (i) The output must accurately reflect the pose of the target cameras, and (ii), when several views of the same object are generated, they must be multi-view consistent. In order to achieve high 3D accuracy and consistency, recent and concurrent methods [42, 44, 63, 83] use a 3D representation of the reconstructed object, usually on top of a pre-trained 2D generative model [57]. It works well but adds complexity.
@@
In this paper, we introduce Free3D, a simple, flexible, and efficient method that can also achieve consistent NVS results without the need to rely on a resource-intensive 3D representation. Zero-1-to-3 [43] is perhaps the best-known example of such a 3D-free NVS system. Like it, Free3D builds upon a pre-trained 2D generative model like Stable Diffusion [57], trained on a large-scale dataset (over 5 billion images [62]), as a data prior. The prior knowledge contained in such a 2D generator is extremely important to be able to ‘guess’ plausible novel views of arbitrary new objects, which is inherently highly ambiguous. However, we show empirically that Zero-1-to-3, has, in practice, poor camera pose control, and, when tasked with generating multiple views, not very consistent. The latter is unavoidable in their design because each view is sampled independently from scratch and, owing to the ambiguity of reconstruction, there is no reason why compatible images would be generated each time. We refer readers to the supplemental videos for a comprehensive comparison.
@@
To mitigate these issues, we first show that much better camera control can be achieved by switching to a different representation of camera pose. Specifically, we introduce a ray conditioning normalization (RCN) (Fig. 2 (b)), which tells each pixel its viewing direction. This is a distributed representation, which should be contrasted to the concentrated camera representation used in [43]. They pass the camera elevation, azimuth, and distance as language-like tokens that may be difficult for the network to interpret and utilize [61]. In contrast, with our RCN layer, we show how to effectively incorporate this per-pixel ray information in an existing text-to-image diffusion model, which empirically leads to significantly more accurate NVS in our experiments (Tabs. 1 and 2). It is also reminiscent of the design of methods like NeRF [48] and LFNs [66], which also work by processing individual rays.
@@
While RCN leads to more accurate camera control, it cannot improve multi-view consistency by itself. To improve the latter, we introduce an additional pseudo-3D cross-view attention module (Fig. 2(c)), inspired by video diffusion models [64, 30, 8, 26, 80], that fuses information across all views instead of processing each view independently. Furthermore, we use multi-view noise sharing when generating the different views of the object, which further enhances consistency due to the continuity of the denoising function, reducing aleatoric variations from view to view.
@@
We benchmark Free3D against recent and concurrent state-of-the-art methods [43, 44, 77, 17] on open-set NVS. Although the model is trained on only one dataset, it generalizes well to all recent NVS benchmark datasets, including Objaverse [18], OmniObject3D [81], and Google Scanned Object (GSO[20]. A thorough experimental assessment shows that our approach consistently outperforms existing methods, both quantitatively and qualitatively.
@@
To summarise, with Free3D we make the following contributions: (i) We introduce the ray conditioning normalization (RCN) layer and show that representing the camera by utilizing a combination of distributed ray conditioning and concentrated pose tokens significantly improves pose accuracy in NVS. (ii) We show that a small multi-view attention module can improve multi-view consistency by exchanging information between views at a low cost. (iii) We find that multi-view noise sharing between different views further improves consistency. (iv) We demonstrate empirically that Free3D achieves consistent NVS without needing a 3D representation and outperforms the existing state-of-the-art models on both pose accuracy and view consistency.
@@ 2 Related Work
Early NVS works relied on epipolar geometry to interpolate between different views of the same scene [14, 16]. A recent breakthrough was to represent 3D scenes as implicit neural fields, as proposed by SRN [65], DeepSDF [51], NeRF [48] and LFN [66], and further improved in follow-ups [87, 5, 6, 24, 11, 49, 38]. Even so, data efficiency, generalizability, and robustness remain a limitation: such systems require multiple views to learn the 3D representation from scratch of every single scene.
To bypass the need for multiple input views, DreamFusion [52] proposed to distill 3D models from a large-scale pre-trained 2D diffusion model [58]. RealFusion [46] extended the latter to single-view image reconstruction by adding the input image as a constraint during distillation. Several follow-up works [70, 54, 61, 53] achieved further improvements to the resolution and quality of the resulting 3D assets. Although these models are open-ended [46], they require lengthy per-scene optimisation.
Figure 2: The overall pipeline of our Free3D. (a) Given a single source input image, the proposed architecture jointly predicts multiple target views, instead of processing them independently. To achieve a consistent novel view synthesis without the need for 3D representation, (b) we first propose a novel ray conditional normalization (RCN) layer, which uses a per-pixel oriented camera ray to module the latent features, enabling the model’s ability to capture more precise viewpoints. (c) A memory-friendly pseudo-3D cross-attention module is introduced to efficiently bridge information across multiple generated views. Note that, here the similarity score is only calculated across multiple views in temporal instead of spatial, resulting in a minimal computational and memory cost.
@@
Inspired by the success of 2D generation [39, 25, 67], some work [71, 50, 78, 15] built the autoencoder architecture for NVS. Driven by the effectiveness of light field rendering (LFR) [2, 1], other generalizable NVS approaches [68, 60, 59] query a network for colour of different rays. 3DiM [76] then introduced diffusion models into NVS. This is followed by RenderDiffusion [3], HoloDiffusion [36], SparseFusion [90], GeNVS [9], Viewset Diffusion [69] and LFD [82]. However, these methods train their models from scratch using pure 3D data. Because such datasets are not sufficient for open-set generalization, they are limited to one or few categories. Here we start from an open-ended 2D image generator instead.
@@
In order to operate in an open-set setting, Zero-1-to-3 [43] built on a 2D image generator trained on a large-scale image dataset [62] and fine-tune it on the Objaverse dataset [18]. While it has good generalizability, Zero-1-to-3 fails to achieve high pose accuracy, and their reconstructions are not very consistent across views. To mitigate these issues, other concurrent approaches either integrate 3D representations into the network [83, 44, 35] or train another auxiliary 3D network [42, 77, 63, 41]. These methods can output high-quality target views, but they are computationally expensive and use significant memory for training. More importantly, these methods do not address the issue of pose representation, which we find to be a key bottleneck in NVS. In contrast, our method is 3D-free and achieves comparable or better NVS quality, due to ray conditioning normalisation, multi-view attention, and multi-view noise sharing. A more recent concurrent work is Stable Video Diffusion [7], which also synthesizes multi-view videos. However, they are trained with fixed frames with limited elevation angles. In contrast, our Free3D is trained to handle arbitrary camera viewpoints, and ideally generates ${}^{\circ}$ spectrum, with flexible-frames for any elevation angles.
@@ 3 Method
Our goal is to learn a model $\Phi$ that, given as input a source image $x^{\text{src}}$ and a sequence of camera poses $\mathcal{P}=\{\mathbf{P}^{i}\}^{N}_{i=1}$, $\mathbf{P}^{i}=(\mathbf{K}^{i},\mathbf{R}^{i},\mathbf{T}^{i})$ comprising intrinsics $\mathbf{K}^{i}$, rotation $\mathbf{R}^{i}$ and translation $\mathbf{T}^{i}$, synthesizes corresponding novel views $\{x^{i}\}_{i=1}^{N}$ which are accurate and consistent without relying on an explicit 3D representation. We approach this problem by generating all the views together, conditioned on the source image $x^{\text{src}}$ and a series of camera parameters. In this way, the network has a chance to produce several consistent views together. The broad motivation here is that the 1D sequential representation has been used extensively in the 2D generation [74, 73, 75, 56, 13, 23, 89]. Analogously, it should be able to produce visually consistent video representing the underlying 3D structure using 2D sequential representation, i.e. multi-frames [7, 4]. However, we tackle it in a bidirectional manner [40, 22, 10, 89, 86], instead of sequentially auto-regressive generation.
To realize this goal, we must tackle two challenges: (i) ensuring that the model accurately captures the target view and (ii) ensuring that different views of the object are consistent in geometry and appearance. To achieve these, our framework, illustrated in Fig. 2, extends a 2D generator by injecting at each layer a ray-conditional normalization (RCN) layer (Sec. 3.1) as well as pseudo-3D attention layer (Sec. 3.2). It also utilizes multi-view noise sharing (Sec. 3.2). The former captures pose more accurately, whereas the last two improve consistency. Although several prior NVS works have started from 2D generators and also focused on improving multi-view consistency, with RCN we are the first (to the best of our knowledge) to use ray conditioning in this setting and obtain high pose accuracy without a 3D representation or an additional 3D network.
@@ 3.1 Ray Conditioning Normalization (RCN)
Ray conditioning was originally proposed for NVS by [66]. However, it has not been used for NVS models that build on 2D generators like Zero-1-to-3 [43] — in practice, doing so is essential for generalization due to the lack of very large-scale 3D data. Here, we thus develop a method to extend a pretrained 2D generator with ray conditioning.
@@
Given a target view $\mathbf{P}=(\mathbf{K},\mathbf{R},\mathbf{T})$, for each pixel $(u,v)$ in the image, we define the Plücker coordinates $\bm{r}_{uv}=\phi(\bm{o},\bm{d}_{uv})=(\bm{o}\times\bm{d}_{uv},\bm{d}_{uv})\in\mathbb{R}^{6}$, of the ray going from the camera center $\bm{o}\in\mathbb{R}^{3}$ through the pixel, while $\bm{d}_{uv}=\mathbf{R}^{\top}(\mathbf{K}^{-1}(u,v,1)^{\top}-\mathbf{T})\in\mathbb{R}^{3}$ is the ray direction. This encoding was originally introduced by LFN [66]. It is invariant to shifting the camera along the ray, meaning that $\phi(\bm{o}+\lambda\bm{d},\bm{d})=((\bm{o}+\lambda\bm{d})\times\bm{d}_{uv},\bm{d}_{uv})=(\bm{o}\times\bm{d}_{uv},\bm{d}_{uv})=\phi(\bm{o},\bm{d})$, which matches the fact that light propagates in straight lines.
@@
  • Concatenation. Following Zero-1-to-3 [43], a natural choice is to concatenate the noised target $z_{t}^{\text{tgt}}$, original source embedding $z^{\text{src}}$, and ray conditioning at the input. Then, the input is now $(z_{t}^{\text{tgt}},z^{\text{src}},\bm{r})$ instead of $z_{t}^{\text{tgt}}$ alone.
  • Multi-scales concatenation. We further consider concatenating ray embeddings $\bm{r}$ to each intermediate layer in the UNet $\epsilon_{\theta}$. Note that each layer operates at a different resolution, so this amounts to injecting the information $\bm{r}$ at different scales.
@@
    where $(\gamma,\beta)=\operatorname{MLP}^{\text{mod}}(\bm{r})$ are the learned scale and shift from ray conditioning $\bm{r}$. Such modulation is applied to each sub-module of UNet $\epsilon_{\theta}$ (Fig. 2 (a) & (b)).
Interestingly, while RCN works best, we show empirically (Tab. 3) that all such architectures lead to strong improvements. This confirms the importance of ray conditioning for the significant improvements, with the novel RCN emerging as the frontrunner in this progression.
@@ Discussion.
RCN differs significantly from Zero-1-to-3 [43] and follow-ups [42, 44, 83, 77, 33] which coded $\mathbf{P}$ as global tokens. Ray-Cond [12] and LFD [82] also applied ray conditioning for NVS in GAN and Diffusion, respectively, but just concatenated it as additional channels. Besides, they are only trained for category-specific NVS. Similar to ray-casting [45, 48], RCN considers one ray per pixel. However, it eschews the computation of hundreds of samples per ray. Because of this, Free3D dramatically reduces the rendering time and memory consumption compared to the concurrent works of [44, 42, 83].

### v2

@@ Abstract
We introduce Free3D, a simple accurate method for monocular open-set novel view synthesis (NVS). Similar to Zero-1-to-3, we start from a pre-trained 2D image generator for generalization, and fine-tune it for NVS. Compared to other works that took a similar approach, we obtain significant improvements without resorting to an explicit 3D representation, which is slow and memory-consuming, and without training an additional network for 3D reconstruction. Our key contribution is to improve the way the target camera pose is encoded in the network, which we do by introducing a new ray conditioning normalization (RCN) layer. The latter injects pose information in the underlying 2D image generator by telling each pixel its viewing direction. We further improve multi-view consistency by using light-weight multi-view attention layers and by sharing generation noise between the different views. We train Free3D on the Objaverse dataset and demonstrate excellent generalization to new categories in new datasets, including OmniObject3D and GSO. The project page is available at https://chuanxiaz.com/free3d/.
@@
Figure 1: Given a single input view, Free3D synthesizes consistent $360^{\circ}$ views accurately without using an explicit 3D representation. Trained on Objaverse only, it generalizes well to new datasets and categories.
@@ 1 Introduction
Novel view synthesis (NVS) has seen significant recent progress [4, 44, 5, 9, 33], in part due to new neural representations like NeRF [43]. However, many such NVS methods require to optimize a 3D model from scratch for each scene, and require dozens of input views to work well; they are thus impractical in many applications. This has motivated authors to apply generative models [34, 20, 61] to the NVS task, eschewing the need for an explicit 3D model and enabling NVS from a single image [45, 29, 78, 27, 68, 42, 63]. Most of these learn 3D priors that are applicable to an entire object category, or even to unstructured collections of objects [45, 29, 78, 27, 68, 42, 63]. In this work, we aim at improving the quality of these approaches while also generalizing further their applicability by considering an open-set setting, where at test time one is given not only new object instances and categories, but also new datasets.
@@
There are two quality targets for NVS: (i) The output must accurately reflect the pose of the target cameras, and (ii), when several views of the same object are generated, they must be mutually consistent. In order to achieve these goals, recent methods [37, 39, 57, 76, 75] build a 3D representation of the object or scene, often combined with a pre-trained 2D generative model [51]. Using a 2D model with a 3D representation works well but adds complexity.
@@
In this paper, we introduce Free3D, a simple and efficient method that can also achieve consistent NVS results without the need to rely on an explicit 3D representation. Zero-1-to-3 [38] and its precursor 3DiM [68] are perhaps the best-known examples of such a 3D-free NVS system. Like Zero-1-to-3, Free3D builds upon a pre-trained 2D generative model like Stable Diffusion [51], trained on a large-scale dataset (i.e. LAION 5B [56]), as a data prior. The prior knowledge contained in such a 2D generator is extremely important to be able to ‘guess’ plausible novel views of open-set objects, which is inherently highly ambiguous. However, we show empirically that Zero-1-to-3 has, in practice, poor camera pose control, and, when tasked with generating multiple views, not very consistent. The latter is unavoidable in their design because each view is sampled independently from scratch and, owing to the ambiguity of reconstruction, there is no reason why compatible images would be generated each time.
@@
To mitigate these issues, we first show that better camera control can be achieved by switching to a different representation of camera pose. Specifically, we introduce a ray conditioning normalization (RCN) layer (Fig. 2 (b)) which tells each pixel its viewing direction. This is a distributed representation of the camera, which should be contrasted to the concentrated camera representation used in [38], where the camera pose is processed as language-like tokens that may be difficult for the network to interpret and utilize [55]. In contrast, with our RCN layer, we show how to effectively incorporate this per-pixel ray information in an existing text-to-image diffusion model, which empirically leads to significantly more accurate NVS in our experiments (Tabs. 1 and 2). RCN is inspired by the design of methods like NeRF [43], LFNs [60] and 3DiM [68], which also work by modelling individual rays.
@@
While RCN allows to control the camera more accurately, it does not improve multi-view consistency. For the latter, we introduce a pseudo-3D cross-view attention module (Fig. 2 (c)) which is inspired by video diffusion models [58, 25, 7, 21, 72, 6]. This layer fuses information across all views instead of processing them independently. Furthermore, we use multi-view noise sharing when generating the different views of the object, which further enhances consistency due to the continuity of the denoising function, reducing aleatoric variations between the views.
@@
We benchmark Free3D against state-of-the-art methods [38, 39, 69, 14] on open-set NVS. Although the model is trained on only one dataset, it generalizes well to all recent NVS benchmark datasets, including Objaverse [15], OmniObject3D [73], and Google Scanned Object (GSO[17]. A thorough experimental assessment shows that our approach consistently outperforms existing methods, both quantitatively and qualitatively.
@@
To summarise, with Free3D we make the following contributions: (i) We introduce the ray conditioning normalization (RCN) layer and show that representing the camera by utilizing a combination of distributed ray conditioning and concentrated pose tokens significantly improves pose accuracy in NVS. (ii) We show that a small multi-view attention module is sufficient to improve multi-view consistency by exchanging information between views at a low cost. (iii) We find that multi-view noise sharing further improves consistency. (iv) We demonstrate empirically that Free3D achieves consistent NVS without needing a 3D representation and outperforms the existing state-of-the-art models on both pose accuracy and view consistency.
@@ 2 Related Work
Figure 2: The overall pipeline of our Free3D. (a) Given a single input image, the Free3D jointly predicts multiple target views, instead of processing them independently. (b) We propose a novel ray conditional normalization (RCN) layer, which uses a per-pixel oriented camera ray to module the latent features, enabling the model’s ability to capture more precise viewpoints. (c) A memory-friendly pseudo-3D cross-attention module is introduced to efficiently bridge information across multiple generated views.
@@
Early NVS works relied on epipolar geometry to interpolate between different views of the same scene [11, 13]. A recent breakthrough was to represent 3D scenes as implicit neural fields, as proposed by SRN [59], DeepSDF [46], NeRF [43] and LFN [60], and further improved in follow-ups [79, 4, 5, 19, 9, 44, 33]. Even so, data efficiency, generalizability, and robustness remain a limitation: such systems require multiple views to learn the 3D representation from scratch for every scene.
To bypass the need for multiple input views, DreamFusion [47] proposes to distill 3D models from a large-scale pre-trained 2D diffusion model [52]. RealFusion [41] extends the latter to single-view image reconstruction by adding the input image as a constraint during distillation. Several follow-up works [64, 49, 55, 48] further improve the resolution and quality of the resulting 3D assets. Although these models are open-ended [41], they still require lengthy per-scene optimisation.
@@
Some recent work [65, 45, 70, 12] built the autoencoder architecture for NVS. Driven by the effectiveness of light field rendering [2, 1], other NVS approaches [62, 54, 53] query a network for colour of different rays. 3DiM [68] then introduced diffusion models into NVS, an approach also taken by RenderDiffusion [3], HoloDiffusion [31], SparseFusion [82], GeNVS [8], Viewset Diffusion [63] and LFD [74]. However, these methods train their models from scratch using pure 3D data, which are too small to afford open-set generalization, and are thus limited to one or few categories.
@@
To operate in an open-set setting, Zero-1-to-3 [38] builds on a large pre-trained 2D image generator and fine-tune it on the Objaverse dataset [15]. While it has good generalizability, it fails to achieve high pose accuracy, and its reconstructions are inconsistent across views. To mitigate these issues, other concurrent approaches either integrate 3D representations into the network [76, 39, 30] or train another 3D network [37, 69, 57, 36, 75]. These methods can output high-quality target views, but are computationally expensive. More importantly, these methods do not address the issue of pose representation, which we find to be a key bottleneck in NVS. In contrast, our method is 3D-free and achieves comparable or better NVS quality, due to ray conditioning normalisation, multi-view attention, and multi-view noise sharing. A more recent concurrent work is SVD [6], which synthesizes multi-view videos, but is trained with fixed camera poses. In contrast, Free3D is trained to generate arbitrary 360 viewpoints. The follow-up work SV3D [67] builds on a similar concept.
@@ 3 Method
Our goal is to learn a model $\Phi$ that, given as input an image $x^{\text{src}}$ and a sequence of camera poses $\mathcal{P}=\{\mathbf{P}^{i}\}^{N}_{i=1}$, synthesizes corresponding novel views $\{x^{i}\}_{i=1}^{N}$ which are accurate and consistent without relying on an explicit 3D representation. We generate all the views together, conditioned on the given inputs, so that the network has a chance to produce several consistent views together.
Specifically, we address two important challenges: (i) ensuring that the model accurately captures the pose of the target view and (ii) ensuring the different views consistent in terms of geometry and appearance. To achieve these, our framework, illustrated in Fig. 2, extends a 2D generator by injecting at each layer a ray-conditional normalization layer (Sec. 3.1) and a pseudo-3D attention layer (Sec. 3.2). It also utilizes multi-view noise sharing (Sec. 3.2). The former captures pose more accurately, whereas the last two improve multi-view consistency. Although 3DiM [68] also passes ray information to the network, the key difference is that we show how to inject ray conditioning in a pre-trained 2D generator. The benefit is to marry generalization with high pose accuracy without an explicit 3D representation or an additional 3D reconstruction network.
@@ 3.1 Ray Conditioning Normalization (RCN)
Given a target view $\mathbf{P}=(\mathbf{K},\mathbf{R},\mathbf{T})$, for each pixel $(u,v)$ in the image, we define the Plücker coordinates $\bm{r}_{uv}=\phi(\bm{o},\bm{d}_{uv})=(\bm{o}\times\bm{d}_{uv},\bm{d}_{uv})\in\mathbb{R}^{6}$ of the ray going from the camera center $\bm{o}\in\mathbb{R}^{3}$ through the pixel, where $\bm{d}_{uv}=\mathbf{R}^{\top}(\mathbf{K}^{-1}(u,v,1)^{\top}-\mathbf{T})\in\mathbb{R}^{3}$ is the ray direction and $(\mathbf{K},\mathbf{R},\mathbf{T})$ are the camera’s calibration, rotation and translation parameters. This encoding was originally introduced by LFN [60]. It is invariant to shifting the camera along the ray, meaning that $\phi(\bm{o}+\lambda\bm{d},\bm{d})=((\bm{o}+\lambda\bm{d})\times\bm{d}_{uv},\bm{d}_{uv})=(\bm{o}\times\bm{d}_{uv},\bm{d}_{uv})=\phi(\bm{o},\bm{d})$, which matches the fact that light propagates in straight lines.
@@
  • Concatenation. Following Zero-1-to-3 [38], a natural choice is to concatenate the noised target $z_{t}^{\text{tgt}}$, original source embedding $z^{\text{src}}$, and ray conditioning at the input, which is then $(z_{t}^{\text{tgt}},z^{\text{src}},\bm{r})$ instead of $z_{t}^{\text{tgt}}$ alone.
  • Multi-scales concatenation. We further consider concatenating the ray embeddings $\bm{r}$ to each intermediate layer in the UNet $\epsilon_{\theta}$. Note that each layer operates at a different resolution, so this amounts to injecting the information $\bm{r}$ at different scales.
@@
    where $(\gamma,\beta)=\operatorname{MLP}^{\text{mod}}(\bm{r})$ are scale and shift parameters predicted from the ray embeddings $\bm{r}$. This is applied to each sub-module of the UNet $\epsilon_{\theta}$ (Fig. 2 (a) & (b)).
Interestingly, while RCN works best, we show empirically that all such architectures lead to strong improvements (Tab. 3), which confirms that the key factor in better camera control is to inject ray information in the network.
@@ Discussion.
RCN differs significantly from Zero-1-to-3 [38] and follow-ups [37, 39, 76, 69, 28, 67] which encode the camera pose $\mathbf{P}$ as global tokens. Ray-Cond [10] and LFD [74] also applied ray conditioning for NVS in GAN and Diffusion, respectively, but they only concatenate it as additional channels, and consider category-specific NVS. Similar to ray-casting [40, 43], RCN considers one ray per pixel, but, differently from the, does not evaluate hundreds of samples per ray. Because of this, Free3D dramatically reduces the rendering time and memory consumption compared to concurrent works like [39, 37, 76].